1) In a certain class of students, there are 11 boys from Wilmette, 6 girls from Winnetka, 7 1) girls from Wilmette, 4 boys from Glencoe, 2 boys from Winnetka and 9 girls from Glenoce.
If the teacher calls upon a student to answer a question, what is the probability that the
student will be a boy?
Answer #1

Total number of boys = 11 + 4 + 2 = 17

Total number of girls = 6 + 7 + 9 = 22

Total number of students = 17 + 22 = 39

P(the student will be a boy) = Number of boys / Total number of students

= 17/39

= 0.4359